OpenStudy (anonymous):

I need help resolving the equation by the method of factorisation: 3x^2+5x + 2 = 0 I thought it was x(2x+4) 1(x+1) x=-4 and x=-1 But it gives me 2x^2 + 5x + 1 Help? :)

OpenStudy (anonymous):

(x+1)(3x+2)

OpenStudy (anonymous):

(3x + 2 )(x + 1 )

OpenStudy (anonymous):

@serrico do u know how the factorisation method works???

OpenStudy (anonymous):

Sorry it was meant to be 2x^2+5x + 2 = 0 @Harkirat, I do is just this question, I have to multiply 2x x 2 = 4 and two numbers that add up to 5 is 4 and 1

OpenStudy (anonymous):

correct u hv to fist take the product of the coefficient of x^2 and the constant here it is 2 x 2 = 4 Now u hv to split 4 into two factors so that their sum is 5 u hv rightly chose 1 and 4 because 1x4 = 4 and 1+5 = 5

OpenStudy (anonymous):

2x^2 + 5x + 2 = 0 2x^2 + 4x + 1x + 2 = 0 2x(x + 2) +1(x + 2) = 0 here u hv to ensure that terms inside bracket r same and do not have any common factors) so (2x + 1)(x+2) = 0 if 2x + 1 = 0 ==> x = -1/2 if x + 2 = 0 ==> x = -2 so yr roots are x = -1/2 , -2

OpenStudy (anonymous):

to show u your mistake, I hv copied a part of yr answer you wrote above "I thought it was x(2x+4) 1(x+1) <------ Mistake here. In first bracket 2 is common in 2x and 4 and so has to be taken out

OpenStudy (anonymous):

is it clear now????????? if so, pls click GOOD ANSWER button for me ☺

OpenStudy (anonymous):

Thank you very much, don't worry I was going to do that I was just waiting until you finished typing and read what you wrote and I understood very clearly. Thank you Harkirat, just what I needed ;)

OpenStudy (anonymous):

u r welcome :)